### v1

# Last Iterate Convergence of Incremental Methods and Applications in Continual Learning

Abstract
Incremental gradient methods and incremental proximal methods are a fundamental class of optimization algorithms used for solving finite sum problems, broadly studied in the literature. Yet, when it comes to their convergence guarantees, nonasymptotic (first-order or proximal) oracle complexity bounds have been obtained fairly recently, almost exclusively applying to the average iterate. Motivated by applications in continual learning, we obtain the first convergence guarantees for the last iterate of both incremental gradient and incremental proximal methods, in general convex smooth (for both) and convex Lipschitz (for the proximal variants) settings. Our oracle complexity bounds for the last iterate nearly match (i.e., match up to a square-root-log or a log factor) the best known oracle complexity bounds for the average iterate, for both classes of methods. We further obtain generalizations of our results to weighted averaging of the iterates with increasing weights, which can be seen as interpolating between the last iterate and the average iterate guarantees. Additionally, we discuss how our results can be generalized to variants of studied incremental methods with permuted ordering of updates. Our results generalize last iterate guarantees for incremental methods compared to state of the art, as such results were previously known only for overparameterized linear models, which correspond to convex quadratic problems with infinitely many solutions.
1 Introduction
We study the last iterate convergence of incremental (gradient and proximal) methods, which apply to problems of the form
$\min_{{\bm{x}}\in\mathbb{R}^{d}}\Big{\{}f({\bm{x}}):=\frac{1}{T}\sum_{t=1}^{T}f_{t}({\bm{x}})\Big{\}}.$ (1.1)
As is standard, we assume that each component function $f_{t}$ is convex and either smooth or Lipschitz-continuous and that a minimizer ${\bm{x}}_{*}\in\operatorname*{arg\,min}_{{\bm{x}}}f({\bm{x}})$ exists.
Incremental methods traverse all the component functions $f_{t}$ in a cyclic manner, updating their iterates by taking either gradient descent steps (in the case of incremental gradient methods) or proximal-point steps (in the case of incremental proximal methods) with respect to the individual component functions $f_{t}$. For a more precise statement of these two classes of methods, see Sections 2 and 3. Same as prior work (Bertsekas et al., 2011; Bertsekas, 2011; Li et al., 2019; Mishchenko et al., 2020; Cai et al., 2023a), we define oracle complexity of these methods as the number of first-order or proximal oracle queries to individual component functions $f_{t}$ required to reach a solution ${\bm{x}}$ with optimality gap $f({\bm{x}})-f({\bm{x}}_{*})\leq\epsilon$ on the worst-case instance from the considered problem class, where $\epsilon>0$ is a given error parameter.
Our main motivation for studying the last iterate convergence of incremental methods comes from applications in continual learning (CL). In particular, CL models a sequential learning setting, where a machine learning model gets updated over time, based on the changing or evolving distribution of the data passed to the learner. A major challenge in such dynamic learning settings is the degradation of model performance on previously seen data, known as the catastrophic forgetting (McCloskey and Cohen, 1989; Goodfellow et al., 2013), which has been well-documented in various empirical studies; see, e.g., recent surveys (De Lange et al., 2021; Parisi et al., 2019). On the theoretical front, however, much is still missing from the understanding of possibilities and limitations related to catastrophic forgetting, with results for basic learning settings being obtained very recently (Evron et al., 2022, 2023; Lin et al., 2023b; Goldfarb and Hand, 2023; Evron et al., 2024; Peng and Risteski, 2022; Peng et al., 2023; Chen et al., 2022; Cao et al., 2022; Balcan et al., 2015).
While there are different learning setting studied under the umbrella of CL, following recent work (Evron et al., 2022, 2023), we focus on the CL settings with repeated replaying of tasks, where the forgetting after $K$ epochs/full passes over $T$ tasks is defined by Doan et al. (2021); Evron et al. (2022, 2023)
$f_{K}({\bm{x}}_{KT}):=\frac{1}{T}\sum_{t=1}^{T}f_{t}({\bm{x}}_{KT}),$ (1.2)
where ${\bm{x}}_{n}\in{\mathbb{R}}^{d}$ is the model parameter vector used at time $n\in\mathbb{N}_{+}$. Such settings arise in applications that naturally undergo cyclic changes in the data/tasks, due to diurnal or seasonal cycles (e.g., in agriculture, forestry, e-commerce, astronomy, etc.).
Observe that (1.2) corresponds to the value of the objective function from (1.1) at the final iterate ${\bm{x}}_{KT}.$ Prior work (Evron et al., 2022) that obtained rigorous bounds for the forgetting (1.2) applied to the problems where each $f_{t}$ is a convex quadratic function minimized by the same ${\bm{x}}_{*}$ such that $f_{t}({\bm{x}}_{*})=f({\bm{x}}_{*})=0.$ By contrast, we consider more general convex functions that are either smooth or Lipschitz continuous, and make no assumption about ${\bm{x}}_{*}$ beyond being a minimizer of the (average) function $f$. Since we are not assuming that $f({\bm{x}}_{*})=0,$ our focus is on bounding the excess forgetting $f_{K}({\bm{x}}_{KT})-f({\bm{x}}_{*})$, which is equivalently the optimality gap for the last iterate in (1.1).
The method considered in Evron et al. (2022) minimized each component function exactly, outputting the solution closest to the previous iterate in each iteration, using implicit regularization properties of SGD. To obtain the results, it was then crucial that the component functions were quadratic (so that there is an explicit, closed-form solution for each subproblem) and that all component functions shared a nonempty set of minimizers with value zero (so that forgetting can be controlled despite aggressive adaption to the current task). Our work using a incremental proximal method instead considers explicit regularization to enforce closeness of models on differing tasks, which can potentially degrade the performance on the current task, but as a tradeoff can control forgetting and it addresses a much broader class of loss functions.
1.1 Contributions
Our main contributions can be summarized as follows.
  • We provide the first oracle complexity guarantees for the last iterate of standard variants of the incremental gradient method with either deterministic or randomly permuted ordering of the updates, applied to convex $L$-smooth objectives. Up to a square-root-logarithm factor, our oracle complexity bounds – which are $\widetilde{\mathcal{O}}\big{(}\frac{TL\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon}+\frac{TL^{1/2}\sigma_{*}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{3/2}}\big{)}$ for the deterministic variant and $\widetilde{\mathcal{O}}\big{(}\frac{TL\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon}+\frac{\sqrt{TL}\sigma_{*}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{3/2}}\big{)}$ for the randomly permuted variant – match the best known oracle complexity bounds for these methods, previously known only for the (uniformly) average iterate (Mishchenko et al., 2020; Nguyen et al., 2021; Cai et al., 2023a). We further extend our results to increasing weighted averaging of iterates, which places more weight on more recent iterates, removing the excess square-root-log factor.
  • We provide the first oracle complexity guarantees for the last iterate of the incremental proximal method, applied to convex and either smooth or Lipschitz-continuous objectives. When each component function is convex and $L$-smooth, we show that the incremental proximal method has the same $\widetilde{\mathcal{O}}\big{(}\frac{TL\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon}+\frac{TL^{1/2}\sigma_{*}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{3/2}}\big{)}$ oracle complexity guarantee as the incremental gradient method. When component functions are convex and $G$-Lipschitz, our oracle complexity $\widetilde{\mathcal{O}}\Big{(}\frac{G^{2}T\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{2}}\Big{)}$ matches the best known oracle complexity bounds up to a logarithmic factor, which were previously known only for the (uniformly) average iterate (Bertsekas, 2011; Li et al., 2019). We further argue that our analysis can be extended to admit inexact proximal point evaluations.
  • We initiate the study of the proximal incremental method as a model of CL, corresponding to sequential ridge-regularized model training commonly used in practice. We discuss how the regularization parameter (or the parameter of proximal-point updates) affects the loss on the current learning task – determined by the regularization error – and (excess) catastrophic forgetting – corresponding to the optimality gap of the last iterate of the method.
1.2 Further Related Work
To remedy catastrophic forgetting, various empirical approaches have been proposed, corresponding to three main categories: (i) memory-based approaches, which store samples from previous tasks and reuse those data for training on the current task (Robins, 1995; Lopez-Paz and Ranzato, 2017; Rolnick et al., 2019); (ii) expansion-based approaches, which progressively expand the model size to accommodate new tasks (Rusu et al., 2016; Yoon et al., 2017); and (iii) regularization-based approaches, which regularize the loss of the current task to ensure the new model parameter vector is close to the old ones (Kirkpatrick et al., 2017); for a more complete survey, we refer to De Lange et al. (2021). Our focus in this work is on (i) and (ii).
On the theoretical front, the majority of existing work for the settings we consider has studied catastrophic forgetting for overparameterized linear models (Evron et al., 2022, 2023; Lin et al., 2023b; Goldfarb and Hand, 2023; Evron et al., 2024; Swartworth et al., 2024) or under the neural tangent kernel regime (Bennani et al., 2020; Doan et al., 2021). Another related line of work studied the task similarity in CL under the teacher-student learning framework (Asanuma et al., 2021; Lee et al., 2021). A recent work showed that an algorithm recursively minimizing the loss of the current task over the solution space of the previous task, named the ideal continual learner, never forgets, assuming all tasks share a common global minimizer (Peng et al., 2023).
When it comes to the cyclic forgetting that this paper focuses on, theoretical results have only been established for linear models (Evron et al., 2022, 2023; Swartworth et al., 2024). In particular, the analysis for linear regression tasks in Evron et al. (2022); Swartworth et al. (2024) crucially relies on the exact minimization of each task using (S)GD to have closed-form updates between tasks, while Evron et al. (2023) uses alternating projections to analyze linear classification tasks. It is unclear how to extend either of these results to general convex loss functions that we consider.
On a technical level, our results are most closely related to the literature on last iterate convergence guarantees for subgradient-based methods and stochastic gradient descent (SGD) and the literature on incremental gradient methods and shuffled SGD. For the former, we draw inspiration from the recent results (Zamani and Glineur, 2023; Liu and Zhou, 2023), which rely on a clever construction of reference points ${\bm{z}}_{k}$ with respect to which a gap quantity $f({\bm{x}}_{k})-f({\bm{z}}_{k-1})$ gets bounded to deduce a bound for the optimality gap $f({\bm{x}}_{k})-f({\bm{x}}_{*})$ of the last iterate ${\bm{x}}_{*}$. For the latter line of work, we generalize the analysis used exclusively for the optimality gap of the (uniformly) average iterate (Mishchenko et al., 2020; Nguyen et al., 2021; Cha et al., 2023; Cai et al., 2023a) to control the gap-like quantities $f({\bm{x}}_{k})-f({\bm{z}}_{k-1})$, which require a more careful argument for controlling all error terms introduced by replacing ${\bm{x}}_{*}$ by ${\bm{z}}_{k-1}$ without introducing spurious unrealistic assumptions about the magnitudes of the component functions’ gradients. We finally note that both these related lines of work concern problems on which progress was made only in the very recent literature. In particular, while the oracle complexity upper bound for the average iterate of SGD in convex Lipschitz-continuous settings has been known for decades and its analysis is routinely taught in optimization and machine learning classes, there were no such results for the last iterate of SGD until 2013 (Shamir and Zhang, 2013) with improvements and generalizations to these results obtained as recently as in the past few months (Liu and Zhou, 2023). On the other hand, obtaining any nonasymptotic convergence guarantees for incremental gradient methods/shuffled SGD had remained open for decades (Bottou, 2009) until a recent line of work (Gürbüzbalaban et al., 2021; Shamir, 2016; Haochen and Sra, 2019; Nagaraj et al., 2019; Ahn et al., 2020; Rajput et al., 2020; Yun et al., 2022; Safran and Shamir, 2020). For smooth convex problems we consider, the convergence results were obtained only in the past few years (Mishchenko et al., 2020; Nguyen et al., 2021; Cha et al., 2023) and improved in Cai et al. (2023a) using a fine-grained analysis inspired by the recent advances in cyclic methods (Song and Diakonikolas, 2023; Cai et al., 2023b; Lin et al., 2023a). However, all those results are for the (uniformly) average iterate, while obtaining convergence results for the last iterate had remained open until this work.
1.3 Notation and Preliminaries
Throughout the paper, we consider the finite $d$-dimensional Euclidean space $(\mathbb{R}^{d},\|\cdot\|)$, where the norm $\|\cdot\|=\sqrt{\left\langle\cdot,\cdot\right\rangle}$ is induced by the standard inner product associated with the space. We let $\mathbb{N}_{+}$ be the set of positive natural numbers and denote $[T]:=\{1,2,\dots,T\}$. Given a proper, convex, lower semicontinuous function $f$, its proximal operator and Moreau envelope are defined by
$\mathrm{prox}_{\eta f}({\bm{x}})=\;$ $\operatorname*{arg\,min}_{{\bm{y}}\in{\mathbb{R}}^{d}}\Big{\{}\frac{1}{2\eta}\|{\bm{y}}-{\bm{x}}\|^{2}+f({\bm{y}})\Big{\}},$
$M_{\eta f}({\bm{x}})=\;$ $\min_{{\bm{y}}\in{\mathbb{R}}^{d}}\Big{\{}\frac{1}{2\eta}\|{\bm{x}}-{\bm{y}}\|^{2}+f({\bm{y}})\Big{\}},$
respectively, for a parameter $\eta>0$. The Moreau envelope can be seen as a smoothing of $f$, which is $\frac{1}{\eta}$-smooth with the gradient $\nabla M_{\eta f}({\bm{x}})=\frac{1}{\eta}({\bm{x}}-\mathrm{prox}_{\eta f}({\bm{x}}))\in\partial f(\mathrm{prox}_{\eta f}({\bm{x}}))$, where $\partial f(\cdot)$ denotes the subdifferential of $f$. To facilitate the discussion on inexact proximal point steps, following Lemaire (1992); Cominetti (1997); Alber et al. (1997); Salzo and Villa (2012); Villa et al. (2013), for a given error parameter $\epsilon\geq 0$ we further let $\partial_{\epsilon}f(\cdot)$ be the $\epsilon$-subdifferential of $f$ defined by (see the the textbook (Hiriart-Urruty and Lemarechal, 2013) for more details)
$\partial_{\epsilon}f({\bm{x}})=\{{\bm{g}}\in\mathbb{R}^{d}\mid f({\bm{y}})\geq f({\bm{x}})+\left\langle{\bm{g}},{\bm{y}}-{\bm{x}}\right\rangle-\epsilon,\forall{\bm{y}}\in\mathbb{R}^{d}\}.$
We make the following assumptions. The first one is made throughout the paper.
Assumption 1.
Each component function $f_{t}$ is convex and there exists a minimizer ${\bm{x}}_{*}\in\operatorname*{arg\,min}_{{\bm{x}}\in{\mathbb{R}}^{d}}f({\bm{x}})$.
By Assumption 1, $f$ is also convex. In nonsmooth settings, we make an additional standard assumption that the component functions are Lipschitz-continuous.
Assumption 2.
Each $f_{t}$ is $G$-Lipschitz $(t\in[T])$, i.e., $|f_{t}({\bm{x}})-f_{t}({\bm{y}})|\leq G\|{\bm{x}}-{\bm{y}}\|$ for any ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d}$; thus $\|g_{t}({\bm{x}})\|\leq G$ for all $g_{t}({\bm{x}})\in\partial f_{t}({\bm{x}})$.
For the smooth settings, we make the following assumption.
Assumption 3.
Each $f_{t}$ is $L$-smooth $(t\in[T])$, i.e., $\|\nabla f_{t}({\bm{x}})-\nabla f_{t}({\bm{y}})\|\leq L\|{\bm{x}}-{\bm{y}}\|$ for any ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d}$.
We remark that 2 and 3 imply that $f$ is also $G$-Lipschitz and $L$-smooth, respectively. The assumption that all the component functions have the same Lipschitz or smoothness constant is possible to relax, in which case the resulting complexity bounds would scale with the average Lipschitz/smoothness constant. This can be done using the techniques from our prior work (Cai et al., 2023a). However, to avoid complicating the analysis and to keep the focus on the intricacies of the last iterate convergence, here we maintain the simpler setup under 2 and 3, and defer generalizations to the setting of distinct component Lipschitz/smoothness constants to later extended version of this paper.
When $f$ is $L$-smooth and convex, we will often make use of the following standard inequality that fully characterizes the class of $L$-smooth convex functions:
$\frac{1}{2L}\|\nabla f({\bm{x}})-\nabla f({\bm{y}})\|^{2}\leq f({\bm{y}})-f({\bm{x}})-\left\langle\nabla f({\bm{x}}),{\bm{y}}-{\bm{x}}\right\rangle,\quad\forall{\bm{x}},{\bm{y}}\in{\mathbb{R}}^{d}.$ (1.3)
Finally, when each $f_{t}$ is smooth, we assume bounded variance at ${\bm{x}}_{*}$, same as prior work (Mishchenko et al., 2020; Nguyen et al., 2021; Tran et al., 2021, 2022; Cai et al., 2023a).
Assumption 4.
The quantity $\sigma_{*}^{2}:=\frac{1}{T}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{x}}_{*})\|^{2}$ is bounded.
2 Incremental Gradient Descent and Shuffled SGD
In this section, we analyze the oracle complexity for the last iterate of incremental gradient descent (IGD), assuming component functions are smooth and convex. In the context of CL, this corresponds to a simplified setup where the learner incrementally performs a single gradient step on each task and cyclically replays the $T$ tasks. Nevertheless, this setup serves as a warmup to the proximal setup we discuss in the next section. Additionally, it is of independent interest as incremental gradient methods are widely used in the optimization and machine learning literature, where despite the lack of prior theoretical justification, it is typically the last iterate that gets output by the algorithm in practice.
We summarize the IGD method in Algorithm 1, assuming the incremental order $1,2,\dots,T$ in each epoch for simplicity and without loss of generality. The oracle complexity for the (uniformly) average iterate of IGD has been shown to be ${\mathcal{O}}(\frac{TL}{\epsilon}+\frac{T\sqrt{L}\sigma_{*}}{\epsilon^{3/2}})$ for an $\epsilon$-optimality gap (Mishchenko et al., 2020; Cai et al., 2023a) under the same assumptions we make here (1, 3, and 4), while, as discussed before, there were no guarantees for either the last iterate or even a weighted average of the iterates. The main result of this section is that the same oracle complexity applies to the last iterate of IGD, up to a square-root-log factor. We then further generalize this result to weighted averages of iterates with increasing weights and to variants with the randomly permuted order of cyclic updates.
Algorithm 1 Incremental Gradient Descent (IGD)
  Input: initial point ${\bm{x}}_{0}$, number of epochs $K$, step size $\{\eta_{k}\}$
  for $k=1:K$ do
     ${\bm{x}}_{k-1,1}={\bm{x}}_{k-1}$
     for $t=1:T$ do
        ${\bm{x}}_{k-1,t+1}={\bm{x}}_{k-1,t}-\eta_{k}\nabla f_{t}({\bm{x}}_{k-1,t})$
     ${\bm{x}}_{k}={\bm{x}}_{k-1,T+1}$
  return  ${\bm{x}}_{K}$
We begin the analysis by deriving a bound on the gap with respect to an arbitrary but fixed reference point ${\bm{z}},$ as summarized in the following lemma. This stands in contrast to arguments deriving bounds on the average iterate, which take ${\bm{z}}={\bm{x}}_{*}.$ While this may seem like a minor difference, it affects the analysis non-trivially: a direct extension of prior arguments would require replacing 4 – which imposes a bound on $\frac{1}{T}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{x}}_{*})\|^{2}$ – with a bound on $\frac{1}{T}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{z}})\|^{2}$ for an arbitrary ${\bm{z}}$, which would be a much stronger requirement.
Lemma 2.1.
Under Assumptions 1 and 3, for any ${\bm{z}}\in\mathbb{R}^{d}$ that is fixed in the $k$-th cycle of Algorithm 1 and any $\alpha>0,\beta>0$ such that $\frac{1}{\alpha}+\frac{1}{\beta}\leq\frac{1}{2}$, if the step sizes $\eta_{k}$ are such that $\eta_{k}\leq\frac{1}{\sqrt{\beta}TL}$, we have for $k\in[K]$
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}\leq\;$ $\eta_{k}^{2}L\sum_{t=1}^{T}\Big{\|}\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{*})\Big{\|}^{2}+\frac{\alpha}{\beta}T\big{(}f({\bm{z}})-f({\bm{x}}_{*})\big{)}$ (2.1)
$+\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}\|^{2}\big{)}.$
Proof.
Since each $f_{t}$ is convex and $L$-smooth, we have for $t\in[T]$ (see (1.3)):
$f_{t}({\bm{x}}_{k})-f_{t}({\bm{x}}_{k-1,t})\leq\;$ $\left\langle\nabla f_{t}({\bm{x}}_{k}),{\bm{x}}_{k}-{\bm{x}}_{k-1,t}\right\rangle-\frac{\|\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t})\|^{2}}{2L},$ (2.2)
$f_{t}({\bm{x}}_{k-1,t})-f_{t}({\bm{z}})\leq\;$ $\left\langle\nabla f_{t}({\bm{x}}_{k-1,t}),{\bm{x}}_{k-1,t}-{\bm{z}}\right\rangle-\frac{\|\nabla f_{t}({\bm{x}}_{k-1,t})-\nabla f_{t}({\bm{z}})\|^{2}}{2L}.$ (2.3)
On the other hand, letting $\Phi_{k-1,t}({\bm{x}}):=\left\langle\nabla f_{t}({\bm{x}}_{k-1,t}),{\bm{x}}\right\rangle+\frac{1}{2\eta_{k}}\|{\bm{x}}_{k-1,t}-{\bm{x}}\|^{2}$, we have $\nabla\Phi_{k-1,t}({\bm{x}}_{k-1,t+1})=\mathbf{0}$ by the update ${\bm{x}}_{k-1,t+1}={\bm{x}}_{k-1,t}-\eta_{k}\nabla f_{t}({\bm{x}}_{k-1,t})$ in Algorithm 1. Observe that $\Phi_{k-1,t}$ is $\frac{1}{\eta_{k}}$-strong convex, and thus we also have
$\Phi_{k-1,t}({\bm{z}})\geq\Phi_{k-1,t}({\bm{x}}_{k-1,t+1})+\frac{1}{2\eta_{k}}\|{\bm{z}}-{\bm{x}}_{k-1,t+1}\|^{2}.$ (2.4)
Summing (2.2) and (2.3) and using (2.4), we conclude that for $t\in[T]$,
$f_{t}({\bm{x}}_{k})-f_{t}({\bm{z}})$
[…]
$-\frac{1}{2L}\|\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t})\|^{2}-\frac{1}{2L}\|\nabla f_{t}({\bm{x}}_{k-1,t})-\nabla f_{t}({\bm{z}})\|^{2}$
$-\frac{1}{2\eta_{k}}\|{\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1,t}\|^{2}+\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1,t}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k-1,t+1}-{\bm{z}}\|^{2}\big{)}.$
Decomposing $\nabla f_{t}({\bm{x}}_{k})=\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t})+\nabla f_{t}({\bm{x}}_{k-1,t})$ and summing over $t\in[T]$ where we recall ${\bm{x}}_{k-1}={\bm{x}}_{k-1,1}$ and ${\bm{x}}_{k}={\bm{x}}_{k-1,T+1}$, we have
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}$
[…]
$+\underbrace{\sum_{t=1}^{T}\left\langle\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t}),{\bm{x}}_{k}-{\bm{x}}_{k-1,t}\right\rangle}_{{\mathcal{T}}_{2}}+\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}\|^{2}\big{)}$
$-\frac{1}{2L}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t})\|^{2}-\frac{1}{2L}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{x}}_{k-1,t})-\nabla f_{t}({\bm{z}})\|^{2}.$
For the term ${\mathcal{T}}_{1}$, we recall that by the IGD update, $\nabla f_{t}({\bm{x}}_{k-1,t})=-\frac{1}{\eta_{k}}({\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1,t})$ and ${\bm{x}}_{k}-{\bm{x}}_{k-1,t+1}=\sum_{s=t+1}^{T}({\bm{x}}_{k-1,s+1}-{\bm{x}}_{k-1,s})$, and thus we have
${\mathcal{T}}_{1}=\;$ $-\frac{1}{\eta_{k}}\sum_{t=1}^{T-1}\sum_{s=t+1}^{T}\left\langle{\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1,t},{\bm{x}}_{k-1,s+1}-{\bm{x}}_{k-1,s}\right\rangle$
$-\frac{1}{2\eta_{k}}\sum_{t=1}^{T}\|{\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1,t}\|^{2}$
=\;= $-\frac{1}{2\eta_{k}}\Big{\|}\sum_{t=1}^{T}({\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1,t})\Big{\|}^{2}=-\frac{1}{2\eta_{k}}\|{\bm{x}}_{k}-{\bm{x}}_{k-1}\|^{2}\leq 0.$
For the term ${\mathcal{T}}_{2}$, noticing that ${\bm{x}}_{k}-{\bm{x}}_{k-1,t}=-\eta_{k}\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{k-1,s})$ and decomposing $\nabla f_{s}({\bm{x}}_{k-1,s})=(\nabla f_{s}({\bm{x}}_{k-1,s})-\nabla f_{s}({\bm{z}}))+(\nabla f_{s}({\bm{z}})-\nabla f_{s}({\bm{x}}_{*}))+\nabla f_{s}({\bm{x}}_{*})$, we use Young’s inequality with parameters $\alpha>0$ and $\beta>0$ to obtain
${\mathcal{T}}_{2}=\;$ $\sum_{t=1}^{T}\Big{\langle}\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t}),-\eta_{k}\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{k-1,s})\Big{\rangle}$
[…]
$+\frac{\alpha\eta_{k}^{2}L}{2}\sum_{t=1}^{T}\Big{\|}\sum_{s=t}^{T}\big{(}\nabla f_{s}({\bm{z}})-\nabla f_{s}({\bm{x}}_{*})\big{)}\Big{\|}^{2}+\eta_{k}^{2}L\sum_{t=1}^{T}\Big{\|}\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{*})\Big{\|}^{2}$
$+\frac{\beta\eta_{k}^{2}L}{2}\sum_{t=1}^{T}\Big{\|}\sum_{s=t}^{T}\big{(}\nabla f_{s}({\bm{x}}_{k-1,s})-\nabla f_{s}({\bm{z}})\big{)}\Big{\|}^{2}.$
Further using that $\|\sum_{i=1}^{n}{\bm{x}}_{i}\|^{2}\leq n\sum_{i=1}^{n}\|{\bm{x}}_{i}\|^{2}$ and combining the above bounds on ${\mathcal{T}}_{1}$ and ${\mathcal{T}}_{2}$, we obtain
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}\leq\;$ $\frac{1}{2L}\Big{(}\frac{1}{\alpha}+\frac{1}{\beta}-\frac{1}{2}\Big{)}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t})\|^{2}$
$+\Big{(}\frac{\beta\eta_{k}^{2}T^{2}L}{2}-\frac{1}{2L}\Big{)}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{x}}_{k-1,t})-\nabla f_{t}({\bm{z}})\|^{2}$
$+\frac{\alpha\eta_{k}^{2}T^{2}L}{2}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{z}})-\nabla f_{t}({\bm{x}}_{*})\|^{2}+\eta_{k}^{2}L\sum_{t=1}^{T}\Big{\|}\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{*})\Big{\|}^{2}$
$+\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}\|^{2}\big{)}.$
To make the first two terms on the right-hand side both nonpositive, we require
$\frac{1}{\alpha}+\frac{1}{\beta}\leq\frac{1}{2},\quad\eta_{k}\leq\frac{1}{\sqrt{\beta}TL}.$
We further bound the term $\sum_{t=1}^{T}\|\nabla f_{t}({\bm{z}})-\nabla f_{t}({\bm{x}}_{*})\|^{2}$ by the smoothness and convexity of each $f_{t}$ as follows:
$\sum_{t=1}^{T}\|\nabla f_{t}({\bm{z}})-\nabla f_{t}({\bm{x}}_{*})\|^{2}\leq\;$ $2L\sum_{t=1}^{T}\big{(}f_{t}({\bm{z}})-f_{t}({\bm{x}}_{*})-\left\langle\nabla f_{t}({\bm{x}}_{*}),{\bm{z}}-{\bm{x}}_{*}\right\rangle\big{)}$
=\;= $2TL\big{(}f({\bm{z}})-f({\bm{x}}_{*})\big{)},$
where in the last equation we used $\nabla f({\bm{x}}_{*})=\mathbf{0}$. Hence, we finally obtain
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}\leq\;$ $\eta_{k}^{2}L\sum_{t=1}^{T}\Big{\|}\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{*})\Big{\|}^{2}+\alpha T^{3}\eta_{k}^{2}L^{2}\big{(}f({\bm{z}})-f({\bm{x}}_{*})\big{)}$
$+\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}\|^{2}\big{)}$
[…]
$+\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}\|^{2}\big{)},$
where we used $\eta_{k}\leq\frac{1}{\sqrt{\beta}TL}$ in the last inequality, thus completing the proof. ∎
Our next step is to specify our choice of the reference point ${\bm{z}}_{k}$ for each epoch $k$. In particular, we consider a sequence of points $\{{\bm{z}}_{k}\}_{k=-1}^{K-1}$ that is recursively defined as a convex combination of ${\bm{x}}_{k}$ and the previous reference point ${\bm{z}}_{k-1}$:
${\bm{z}}_{k}=(1-\lambda_{k}){\bm{x}}_{k}+\lambda_{k}{\bm{z}}_{k-1}$ (2.5)
for $k\geq 0$ with ${\bm{z}}_{-1}={\bm{x}}_{*}$ and $\lambda_{k}\in(0,1)$ to be set later.
Observe that ${\bm{z}}_{k}$ can also be written as a convex combination of the points $\{{\bm{x}}_{j}\}_{j=0}^{k}$ and ${\bm{x}}_{*}$ by unrolling the recursion, i.e.,
${\bm{z}}_{k}=(1-\lambda_{k}){\bm{x}}_{k}+\big{(}\prod_{i=0}^{k}\lambda_{i}\big{)}{\bm{x}}_{*}+\sum_{j=0}^{k-1}\big{(}\prod_{i=j+1}^{k}\lambda_{i}\big{)}(1-\lambda_{j}){\bm{x}}_{j},$ (2.6)
where $(1-\lambda_{k})+\prod_{i=0}^{k}\lambda_{i}+\sum_{j=0}^{k-1}\big{(}\prod_{i=j+1}^{k}\lambda_{i}\big{)}(1-\lambda_{j})=1$. If we set $\lambda_{k}=1$ for all $k$, then we have ${\bm{z}}_{k}={\bm{x}}_{*}$ and recover the bound $f({\bm{x}}_{k})-f({\bm{x}}_{*})$ in Lemma 2.1 which leads to the average iterate guarantee. For general $\{\lambda_{k}\}$, we prove the following lemma to relate the function value gap $f({\bm{x}}_{k})-f({\bm{z}}_{k-1})$ to the optimality gap $f({\bm{x}}_{k})-f({\bm{x}}_{*})$.
Lemma 2.2.
Let ${\bm{z}}_{k}$ be defined by (2.5) for a given sequence of parameters $\lambda_{k}\in(0,1)$, where $k\geq 0$ and ${\bm{z}}_{-1}={\bm{x}}_{*}.$ Under 1, if there exists a sequence of nonnegative weights $w_{k}$ such that $\lambda_{k}w_{k}\leq w_{k-1}$ for $k\in[K-1]$, then for all $k\in[K]$:
  1. 1.
    $w_{k-1}\big{(}f({\bm{z}}_{k-1})-f({\bm{x}}_{*})\big{)}\leq\sum_{j=0}^{k-1}w_{j}(1-\lambda_{j})\big{(}f({\bm{x}}_{j})-f({\bm{x}}_{*})\big{)};$
  2. 2.
    $w_{k-1}\big{(}f({\bm{x}}_{k})-f({\bm{z}}_{k-1})\big{)}\geq w_{k-1}(f({\bm{x}}_{k})-f({\bm{x}}_{*}))-\sum_{j=0}^{k-1}w_{j}(1-\lambda_{j})(f({\bm{x}}_{j})-f({\bm{x}}_{*}))$.
Proof.
  1. 1.
    Using (2.6) and the convexity of $f$, we have
    $f({\bm{z}}_{k-1})-f({\bm{x}}_{*})$
[…]
    It remains to multiply by $w_{k-1}$ on both sides and notice that by the lemma assumption,
    $w_{k-1}\big{(}\prod_{i=j+1}^{k-1}\lambda_{i}\big{)}(1-\lambda_{j})\leq w_{k-2}\big{(}\prod_{i=j+1}^{k-2}\lambda_{i}\big{)}(1-\lambda_{j})\leq\cdots\leq w_{j}(1-\lambda_{j}).$
  2. 2.
    This follows from the first part of the lemma, by decomposing $f({\bm{x}}_{k})-f({\bm{z}}_{k-1})=f({\bm{x}}_{k})-f({\bm{x}}_{*})-\big{(}f({\bm{z}}_{k-1})-f({\bm{x}}_{*})\big{)}$.
The role of the sequence of weights $\{w_{k}\}$ in Lemma 2.2 is to ensure that we can telescope the terms $\|{\bm{x}}_{k-1}-{\bm{z}}_{k-1}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}_{k-1}\|^{2}$ in (2.1). On the other hand, to succinctly see why such $\{{\bm{z}}_{k}\}$ could lead to the desired last iterate guarantees, we note that the second part of Lemma 2.2 indicates that $w_{k-1}\big{(}f({\bm{x}}_{k})-f({\bm{z}}_{k-1})\big{)}$ intrinsically includes retraction terms of the optimality gaps at the previous iterates. Hence, we can deduct $w_{K-1}(f({\bm{x}}_{K})-f({\bm{x}}_{*}))$ from $\sum_{k=1}^{K}w_{k-1}\big{(}f({\bm{x}}_{k})-f({\bm{z}}_{k-1})\big{)}$ by properly choosing $\lambda_{k}$ and $w_{k}$ to cancel out the optimality gap terms at the intermediate iterates. In this case, the convergence rate for the last iterate is characterized by the growth rate of the sequence $\{w_{k}\}$.
Our choice of the reference points $\{{\bm{z}}_{k}\}$ is inspired by the recent works (Zamani and Glineur, 2023; Liu and Zhou, 2023) on last iterate guarantees for subgradient methods and SGD with $\lambda_{k}=\frac{w_{k-1}}{w_{k}}$. However, their proof techniques are not directly applicable to incremental methods, due to several technical obstacles including additional nontrivial error terms of the form $\frac{\alpha}{\beta}T\big{(}f({\bm{z}}_{k})-f({\bm{x}}_{*})\big{)}$ in (2.1) arising from the incremental gradient steps using reference points other than ${\bm{x}}_{*}$. Such error terms inherently deteriorate the growth rate of $\{w_{k}\}$ and could possibly lead to a worse last iterate rate compared to the rate on the average iterate. In the following theorem, we calibrate such degradation on the last iterate guarantees, and show that with a slightly smaller step size one can still achieve essentially the same rate as for the average iterate.
Theorem 2.3 (IGD).
Under Assumptions 1, 3, and 4 and for parameters $\alpha,\beta$ such that $\frac{1}{\alpha}+\frac{1}{\beta}\leq\frac{1}{2}$, if the step size is fixed and satisfies $\eta_{k}=\eta\leq\frac{1}{\sqrt{\beta}TL}$, the output ${\bm{x}}_{K}$ of Algorithm 1 satisfies
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\mathrm{e}\eta^{2}T^{2}\sigma_{*}^{2}L(1+\beta/\alpha)K^{\frac{\alpha/\beta}{1+\alpha/\beta}}+\frac{\mathrm{e}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{2T\eta K^{\frac{1}{1+\alpha/\beta}}}.$ (2.7)
With $\alpha=4,\beta=4\log K$, there exists a constant step size $\eta$ such that $f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\epsilon$ for $\epsilon>0$ after $\widetilde{\mathcal{O}}\big{(}\frac{TL\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon}+\frac{TL^{1/2}\sigma_{*}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{3/2}}\big{)}$ individual gradient evaluations.
Proof.
Plugging ${\bm{z}}_{k-1}$ defined by (2.5) into (2.1) from Lemma 2.1 and using the inequality $\|\sum_{i=1}^{n}{\bm{x}}_{i}\|^{2}\leq n\sum_{i=1}^{n}\|{\bm{x}}_{i}\|^{2}$ to bound the term $\sum_{t=1}^{T}\|\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{*})\|^{2}$, we obtain
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}}_{k-1})\big{)}\leq\;$ $T^{3}\eta_{k}^{2}\sigma_{*}^{2}L+\frac{\alpha}{\beta}T\big{(}f({\bm{z}}_{k-1})-f({\bm{x}}_{*})\big{)}$
$+\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1}-{\bm{z}}_{k-1}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}_{k-1}\|^{2}\big{)},$
where $\frac{1}{\alpha}+\frac{1}{\beta}\leq\frac{1}{2}$. Multiplying $\eta_{k}w_{k-1}$ on both sides with $w_{k}$ such that $\lambda_{k}w_{k}\leq w_{k-1}$ and noticing that $\|{\bm{x}}_{k-1}-{\bm{z}}_{k-1}\|^{2}\leq\lambda_{k-1}\|{\bm{x}}_{k-1}-{\bm{z}}_{k-2}\|^{2}$ by (2.5), we have
$T\eta_{k}w_{k-1}\big{(}f({\bm{x}}_{k})-f({\bm{z}}_{k-1})\big{)}\leq\;$ $T^{3}\eta_{k}^{3}\sigma_{*}^{2}Lw_{k-1}+\frac{\alpha}{\beta}T\eta_{k}w_{k-1}\big{(}f({\bm{z}}_{k-1})-f({\bm{x}}_{*})\big{)}$
$+\frac{1}{2}\big{(}w_{k-2}\|{\bm{x}}_{k-1}-{\bm{z}}_{k-2}\|^{2}-w_{k-1}\|{\bm{x}}_{k}-{\bm{z}}_{k-1}\|^{2}\big{)}$
We then sum over $k\in[K]$ and use Lemma 2.2 to obtain
$T\sum_{k=1}^{K}\eta_{k}\Big{[}w_{k-1}\big{(}f({\bm{x}}_{k})-f({\bm{x}}_{*})\big{)}-\sum_{j=0}^{k-1}w_{j}(1-\lambda_{j})\big{(}f({\bm{x}}_{j})-f({\bm{x}}_{*})\big{)}\Big{]}$ (2.8)
[…]
where we also use ${\bm{z}}_{-1}={\bm{x}}_{*}$. Unrolling the terms w.r.t. $f({\bm{x}}_{k})-f({\bm{x}}_{*})$ $(k\in[K])$ we get
$\sum_{k=1}^{K}\eta_{k}\Big{[}w_{k-1}\big{(}f({\bm{x}}_{k})-f({\bm{x}}_{*})\big{)}-\sum_{j=0}^{k-1}w_{j}(1-\lambda_{j})\big{(}f({\bm{x}}_{j})-f({\bm{x}}_{*})\big{)}\Big{]}$ (2.9)
=\;= $\eta_{K}w_{K-1}\big{(}f({\bm{x}}_{K})-f({\bm{x}}_{*})\big{)}-w_{0}(1-\lambda_{0})\big{(}f({\bm{x}}_{0})-f({\bm{x}}_{*})\big{)}\sum_{k=1}^{K}\eta_{k}$
$+\sum_{k=1}^{K-1}\Big{(}\eta_{k}w_{k-1}-w_{k}(1-\lambda_{k})\sum_{j=k+1}^{K}\eta_{j}\Big{)}\big{(}f({\bm{x}}_{k})-f({\bm{x}}_{*})\big{)}$
and
$\sum_{k=1}^{K}\eta_{k}\sum_{j=0}^{k-1}w_{j}(1-\lambda_{j})\big{(}f({\bm{x}}_{j})-f({\bm{x}}_{*})\big{)}=\sum_{k=0}^{K-1}\Big{(}w_{k}(1-\lambda_{k})\sum_{j=k+1}^{K}\eta_{j}\Big{)}\big{(}f({\bm{x}}_{k})-f({\bm{x}}_{*})\big{)}.$
Plugging back into (2.8), grouping the like terms, and choosing $\lambda_{0}=1$, we obtain
$T\eta_{K}w_{K-1}\big{(}f({\bm{x}}_{K})-f({\bm{x}}_{*})\big{)}$ (2.10)
$+T\sum_{k=1}^{K-1}\Big{[}\eta_{k}w_{k-1}-\big{(}1+\frac{\alpha}{\beta}\big{)}w_{k}(1-\lambda_{k})\sum_{j=k+1}^{K}\eta_{j}\Big{]}\big{(}f({\bm{x}}_{k})-f({\bm{x}}_{*})\big{)}$
[…]
To obtain the last iterate guarantee, it suffices to choose $w_{k}$ and $\lambda_{k}$ such that
$\lambda_{k}w_{k}\leq\;$ $w_{k-1},\quad 0\leq k\leq K-1,$ (2.11)
$\eta_{k}w_{k-1}-\big{(}1+\frac{\alpha}{\beta}\big{)}w_{k}(1-\lambda_{k})\sum_{j=k+1}^{K}\eta_{j}\geq\;$ $1\displaystyle 0,\quad 1\leq k\leq K-$ 1 . (2.12)
Noticing that (2.12) is equivalent to $\lambda_{k}\geq 1-\frac{\eta_{k}w_{k-1}}{\big{(}1+\frac{\alpha}{\beta}\big{)}w_{k}\sum_{j=k+1}^{K}\eta_{j}}$, to have both inequalities satisfied at the same time, it suffices that
$1-\frac{\eta_{k}w_{k-1}}{\big{(}1+\frac{\alpha}{\beta}\big{)}w_{k}\sum_{j=k+1}^{K}\eta_{j}}\leq\frac{w_{k-1}}{w_{k}}\iff w_{k}\leq\frac{\eta_{k}+\big{(}1+\frac{\alpha}{\beta}\big{)}\sum_{j=k+1}^{K}\eta_{j}}{\big{(}1+\frac{\alpha}{\beta}\big{)}\sum_{j=k+1}^{K}\eta_{j}}w_{k-1}.$
To maximize the growth rate of $\{w_{k}\}$, we let $w_{k}=\frac{\eta_{k}+(1+\frac{\alpha}{\beta})\sum_{j=k+1}^{K}\eta_{j}}{(1+\frac{\alpha}{\beta})\sum_{j=k+1}^{K}\eta_{j}}w_{k-1}$. Without loss of generality, we take $w_{-1}=w_{0}=\prod_{k=1}^{K-1}\frac{(1+\frac{\alpha}{\beta})\sum_{j=k+1}^{K}\eta_{j}}{\eta_{k}+(1+\frac{\alpha}{\beta})\sum_{j=k+1}^{K}\eta_{j}}$, and thus $w_{K-1}=1$. Hence, dividing both sides of (2.10) by $T\eta_{K}w_{K-1}$ and choosing the constant step size $\eta_{k}\equiv\eta$ for all $k\in[K]$, we obtain
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\eta^{2}T^{2}\sigma_{*}^{2}L\sum_{k=1}^{K}w_{k-1}+\frac{w_{-1}}{2\eta T}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}.$ (2.13)
We first bound $w_{-1}=\prod_{k=1}^{K-1}\frac{(1+\frac{\alpha}{\beta})(K-k)}{1+(1+\frac{\alpha}{\beta})(K-k)}=\prod_{k=1}^{K-1}\frac{(1+\frac{\alpha}{\beta})k}{1+(1+\frac{\alpha}{\beta})k}$ with the constant step size. Taking the natural logarithm of $w_{-1}$, we have
$\sum_{k=1}^{K-1}\log\Big{(}1-\frac{1}{1+\big{(}1+\frac{\alpha}{\beta}\big{)}k}\Big{)}\overset{(i)}{\leq}-\sum_{k=1}^{K-1}\frac{1}{1+\big{(}1+\frac{\alpha}{\beta}\big{)}k}\leq-\frac{1}{1+\frac{\alpha}{\beta}}\sum_{k=1}^{K-1}\frac{1}{k+1}.$
where for $(i)$ we use the fact that $\log(1+x)\leq x$ for $x>-1$. Further noticing that $\sum_{k=1}^{K-1}\frac{1}{k+1}=\sum_{k=1}^{K}\frac{1}{k}-1\geq\log(K)+\frac{1}{K}-1$, then we have
$\log(w_{-1})\leq-\frac{1}{1+\frac{\alpha}{\beta}}\log(K)+\frac{1}{1+\frac{\alpha}{\beta}}\iff w_{-1}\leq\frac{\mathrm{e}^{1/(1+\alpha/\beta)}}{K^{\frac{1}{1+\alpha/\beta}}}\leq\frac{\mathrm{e}}{K^{\frac{1}{1+\alpha/\beta}}}.$
On the other hand, we note that $w_{k-1}=\prod_{j=k}^{K-1}\frac{(1+\frac{\alpha}{\beta})(K-j)}{1+(1+\frac{\alpha}{\beta})(K-j)}=\prod_{j=1}^{K-k}\frac{(1+\frac{\alpha}{\beta})j}{1+(1+\frac{\alpha}{\beta})j}$ for $11\leq k\leq K-$ 1 and $w_{K-1}=1$, then we follow the above argument and obtain
$\sum_{k=1}^{K}w_{k-1}=\;$ $1+\sum_{k=1}^{K-1}\prod_{j=1}^{K-k}\frac{(1+\frac{\alpha}{\beta})j}{1+(1+\frac{\alpha}{\beta})j}$
[…]
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\mathrm{e}\eta^{2}T^{2}\sigma_{*}^{2}L(1+\beta/\alpha)K^{\frac{\alpha/\beta}{1+\alpha/\beta}}+\frac{\mathrm{e}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{2T\eta K^{\frac{1}{1+\alpha/\beta}}}.$
To analyze the oracle complexity, we take
$\eta=\min\Big{\{}\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2/3}}{2^{1/3}T\sigma_{*}^{2/3}L^{1/3}K^{1/3}(1+\beta/\alpha)^{1/3}},\,\frac{1}{\sqrt{\beta}TL}\Big{\}}$
and analyze the two possible cases depending on which term in the min is smaller. If the first term in the min is smaller (which we can equivalently think of as $K$ being “large”), we get
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\frac{3.5(1+\beta/\alpha)^{1/3}L^{1/3}\sigma_{*}^{2/3}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{4/3}}{K^{\frac{1}{1+\alpha/\beta}-\frac{1}{3}}}.$
Alternatively, if $\eta=\frac{1}{\sqrt{\beta}TL}\leq\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2/3}}{2^{1/3}T\sigma_{*}^{2/3}L^{1/3}K^{1/3}(1+\beta/\alpha)^{1/3}}$ (which we can think of as having “small” $K$), we obtain
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\frac{1.8(1+\beta/\alpha)^{1/3}L^{1/3}\sigma_{*}^{2/3}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{4/3}}{K^{\frac{1}{1+\alpha/\beta}-\frac{1}{3}}}+\frac{1.4\sqrt{\beta}L\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{K^{\frac{1}{1+\alpha/\beta}}}.$
Hence, combining these two cases, we have
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq$ $\frac{3.5(1+\beta/\alpha)^{1/3}L^{1/3}\sigma_{*}^{2/3}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{4/3}}{K^{\frac{1}{1+\alpha/\beta}-\frac{1}{3}}}+\frac{1.4\sqrt{\beta}L\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{K^{\frac{1}{1+\alpha/\beta}}}.$
In particular, if we choose $\alpha=4$ and $\beta=4\log K$, assuming without loss of generality that $\log K>1$, then we have $K^{\frac{\alpha/\beta}{1+\alpha/\beta}}=K^{\frac{1}{\log K+1}}\leq\mathrm{e}$ and thus
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\frac{9.4L^{\frac{1}{3}}\sigma_{*}^{\frac{2}{3}}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{\frac{4}{3}}(1+\log K)^{\frac{1}{3}}}{K^{\frac{2}{3}}}+\frac{7.4L\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}\sqrt{\log K}}{K}.$
To guarantee $f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\epsilon$ given $\epsilon>0$, the total number of individual gradient evaluations will be
$TK=\widetilde{\mathcal{O}}\Big{(}\frac{TL\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon}+\frac{TL^{1/2}\sigma_{*}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{3/2}}\Big{)}.$
Observe that the error term $f({\bm{z}}_{k-1})-f({\bm{x}}_{*})$ in (2.1) plays the role of slowing the last iterate rate, as is calibrated by its coefficient $\alpha/\beta$ in (2.7). To remedy such degradation compared to the average iterate rate (Mishchenko et al., 2020; Cai et al., 2023a), one natural thought is to make $\alpha/\beta$ sufficiently small. In particular, we choose $\alpha/\beta=1/\log K$ and show that the last iterate rate nearly matches the best known rate on the average iterate, with the trade-off of requiring order-$\frac{1}{\sqrt{\log K}}$ smaller step sizes in comparison with Mishchenko et al. (2020); Cai et al. (2023a). This translates into the oracle complexity that is larger by at most a $\sqrt{\log(1/\epsilon)}$ factor. For most cases of interest, this quantity can be treated as a constant; in particular, for $\epsilon=10^{-8},$ $\sqrt{\log(1/\epsilon)}\approx 4.$
On the other hand, we note that (2.11) and (2.12) hold trivially with $\lambda_{k}\equiv 1$ and constant $w_{k}$, and then (2.10) directly implies the average iterate guarantee, as a sanity check. Also, instead of saturating (2.11) and (2.12) to obtain the last iterate guarantee, one can deduce the convergence rate on the increasing weighted averaging which places more weight on later iterates, as formalized in the following corollary.
Corollary 2.4 (Increasing Weighted Averaging).
Under Assumptions 1, 3, and 4 and for parameters $\alpha,\beta$ such that $\frac{1}{\alpha}+\frac{1}{\beta}\leq\frac{1}{2}$, if the constant step size $\eta\leq\frac{1}{\sqrt{\beta}TL}$, then for any constant $c\in(0,1]$ and increasing sequence $\{w_{k}\}_{k=0}^{K-1}$ with $w_{k}=\frac{(1+\frac{\alpha}{\beta})(K-k)+1-c}{(1+\frac{\alpha}{\beta})(K-k)}w_{k-1}$lgorithm 1 outputs $\hat{\bm{x}}_{K}=\sum_{k=1}^{K}\frac{w_{k-1}{\bm{x}}_{k}}{\sum_{k=1}^{K}w_{k-1}}$ satisfying
$f(\hat{\bm{x}}_{K})-f({\bm{x}}_{*})\leq\frac{T^{2}\sigma_{*}^{2}\eta^{2}L}{c}+\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{2c\eta TK}.$
In particular, there exists a constant step size $\eta$ such that $f(\hat{\bm{x}}_{K})-f({\bm{x}}_{*})\leq\epsilon$ for $\epsilon>0$ after ${\mathcal{O}}\big{(}\frac{TL\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{c\epsilon}+\frac{TL^{1/2}\sigma_{*}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{c^{3/2}\epsilon^{3/2}}\big{)}$ individual gradient evaluations.
Proof.
We follow the proof of Theorem 2.3 up to (2.10) with constant step size $\eta$, then we instead take $\lambda_{k}=w_{k-1}/w_{k}$ and $w_{k}=\frac{(1+\frac{\alpha}{\beta})(K-k)+1-c}{(1+\frac{\alpha}{\beta})(K-k)}w_{k-1}$ to obtain
$c\eta T\sum_{k=1}^{K}w_{k-1}\big{(}f({\bm{x}}_{k})-f({\bm{x}}_{*})\big{)}\leq T^{3}\eta^{3}\sigma_{*}^{2}L\sum_{k=1}^{K}w_{k-1}+\frac{w_{-1}}{2}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}.$
Since $f$ is convex, we have $f(\hat{\bm{x}}_{K})-f({\bm{x}}_{*})\leq\frac{\sum_{k=1}^{K}w_{k-1}(f({\bm{x}}_{k})-f({\bm{x}}_{*}))}{\sum_{k=1}^{K}w_{k-1}}$ where $\hat{\bm{x}}_{K}=\sum_{k=1}^{K}\frac{w_{k-1}{\bm{x}}_{k}}{\sum_{k=1}^{K}w_{k-1}}$ is the increasing weighted averaging of $\{{\bm{x}}_{k}\}_{k=1}^{K}$, thus (cf. (2.13))
$f(\hat{\bm{x}}_{K})-f({\bm{x}}_{*})\leq\frac{T^{2}\sigma_{*}^{2}\eta^{2}L}{c}+\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}w_{-1}}{2c\eta T\sum_{k=1}^{K}w_{k-1}}\leq\frac{T^{2}\sigma_{*}^{2}\eta^{2}L}{c}+\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{2c\eta TK},$
where the last step is due to $\sum_{k=1}^{K}w_{k-1}\geq Kw_{-1}$. Then we follow the proof of Theorem 2.3 and choose $\eta=\min\{\frac{1}{\sqrt{\beta}TL},(\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{2T^{3}\sigma_{*}^{2}LK})^{1/3}\}$ to obtain
$f(\hat{\bm{x}}_{K})-f({\bm{x}}_{*})\leq\frac{\sqrt{\beta}L\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{2cK}+\frac{2^{1/3}L^{1/3}\sigma_{*}^{2/3}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{4/3}}{cK^{2/3}}.$
To guarantee $f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\epsilon$ for $\epsilon>0$, we choose $\beta={\mathcal{O}}(1)$ and the total number of individual gradient evaluations will be
$TK={\mathcal{O}}\Big{(}\frac{TL\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{c\epsilon}+\frac{TL^{1/2}\sigma_{*}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{c^{3/2}\epsilon^{3/2}}\Big{)}.$
We remark that increasing weighted averaging shaves off the (at most) square-root-log term appearing in the last iterate rate above and recovers the best known rate for the average iterate (Mishchenko et al., 2020; Cai et al., 2023a). The parameters $\alpha,\beta$ are included in the weights $w_{k}$ controlling the growth rate of the increasing sequence $\{w_{k}\}$. In particular, when $\alpha/\beta\rightarrow\infty$, increasing weighted averaging reduces to the uniform weighted average.
In the final part of this section, we extend our analysis to handle the case with possible random permutations on the task ordering of each epoch, showing order-$\sqrt{T}$ improvements in complexity if involving randomness. We consider two main permutation strategies of particular interests in the literature on shuffled SGD: (i) random reshuffling (RR): randomly generate permutations at the beginning of each epoch; (ii) shuffle-once (SO): generate a single random permutation at the beginning and use it in all epochs. We state the convergence results with permutations in the following corollary. Our last iterate guarantee nearly matches the best known average iterate convergence rate for shuffled SGD (Mishchenko et al., 2020; Nguyen et al., 2021; Cai et al., 2023a) and the lower bound results on the RR scheme (Cha et al., 2023), with a slightly (order-$(1/\sqrt{\log K})$) smaller step size.
Corollary 2.5 (Shuffled SGD (RR/SO)).
Under Assumptions 13 and 4 and for parameters $\alpha,\beta$ such that $\frac{1}{\alpha}+\frac{1}{\beta}\leq\frac{1}{2}$, if the constant step size $\eta\leq\frac{1}{\sqrt{\beta}TL}$, the output ${\bm{x}}_{K}$ of Algorithm 1 with uniformly random (SO/RR) shuffling satisfies
$\mathbb{E}[f({\bm{x}}_{K})-f({\bm{x}}_{*})]\leq\;$ $\mathrm{e}\eta^{2}T^{2}\sigma_{*}^{2}L(1+\beta/\alpha)K^{\frac{\alpha/\beta}{1+\alpha/\beta}}+\frac{\mathrm{e}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{2T\eta K^{\frac{1}{1+\alpha/\beta}}}.$
With $\alpha=4,\beta=4\log K$, there exists a constant step size $\eta$ such that $\mathbb{E}[f({\bm{x}}_{K})-f({\bm{x}}_{*})]\leq\epsilon$ for $\epsilon>0$ after $\widetilde{\mathcal{O}}\big{(}\frac{TL\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon}+\frac{\sqrt{TL}\sigma_{*}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{3/2}}\big{)}$ individual gradient evaluations.
Proof.
We first introduce a lemma to bound the term $\sum_{t=1}^{T}\|\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{*})\|^{2}$ in (2.1) with random permutations involved. Its proof is provided in Appendix A.
Lemma 2.6.
Under 4 and for Algorithm 1 with uniformly random (SO/RR) shuffling, we have
$\mathbb{E}\Big{[}\sum_{t=1}^{T}\|\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{*})\|^{2}\Big{]}\leq\frac{T(T+1)}{6}\sigma_{*}^{2}.$
We then take expectation w.r.t. all the randomness on both sides of (2.1) and use Lemma 2.6. Then it remains to follow the analysis in the proof of Theorem 2.3. ∎
3 Incremental Proximal Method
In this section, we leverage the proof techniques developed in the previous section and derive the last iterate convergence bound for the incremental proximal method (IP) summarized in Algorithm 2. While the incremental proximal method is a fundamental method broadly studied in the optimization literature and thus the last iterate convergence bounds are of independent interest, our main motivation for considering this problem comes from CL applications, as discussed in the introduction. Thus we begin this section by briefly explaining this connection and reasoning.
The considered setup is motivated by the general $\ell_{2}$-regularized CL setting (Heckel, 2022; Li et al., 2023). In particular, each proximal iteration ${\bm{x}}_{k-1,t+1}=\mathrm{prox}_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})$ can be interpreted as minimizing the $\ell_{2}$ (a.k.a. ridge) regularized loss $f_{t}$ corresponding to the current task $t$, which aligns with the common machine learning practice of using regularization to improve the generalization error and prevent forgetting. When $\eta_{k}\rightarrow\infty$, the proximal point step reduces to the CL setting where the learner exactly minimizes the loss of the current task, i.e., ${\bm{x}}_{k-1,t+1}=\operatorname*{arg\,min}_{{\bm{x}}\in\mathbb{R}^{d}}f_{t}({\bm{x}})$, while the regularization effect vanishes and causes larger forgetting on previous tasks. When $\eta_{k}$ is small, the proximal point step is easier to compute with larger quadratic regularization and prevents deviating from the previous iterate thus causing less forgetting. However, in this case the plasticity of the model may be deteriorated.
We also note that our analysis of the incremental proximal method is related to previous work on cyclic replays for overparameterized linear models with (S)GD (Evron et al., 2022; Swartworth et al., 2024), as (S)GD in this case acts as an implicit regularizer (Gunasekar et al., 2018; Zhang et al., 2021) (whereas proximal point update acts as an explicit regularizer). The two lines of work are not directly comparable: the prior work (Evron et al., 2022; Swartworth et al., 2024) considers exact minimization of component/task loss function and bounds the forgetting, but only addresses convex quadratics where the component loss functions have a nonempty intersecting set of minima (which implies $\sigma_{*}=0$ in 4). On the other hand, our work addresses much more general smooth convex functions (not necessarily quadratic) and does not require $\sigma_{*}=0,$ but instead relies on sufficiently large regularization.
Algorithm 2 Incremental Proximal Method (IP)
  Input: initial point ${\bm{x}}_{0}$, number of epochs $K$, step size $\{\eta_{k}\}$
  for $k=1:K$ do
     ${\bm{x}}_{k-1,1}={\bm{x}}_{k-1}$
     for $t=1:T$ do
        ${\bm{x}}_{k-1,t+1}=\mathrm{prox}_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})=\operatorname*{arg\,min}_{{\bm{x}}\in\mathbb{R}^{d}}\big{\{}\frac{1}{2\eta_{k}}\|{\bm{x}}-{\bm{x}}_{k-1,t}\|^{2}+f_{t}({\bm{x}})\big{\}}$
     ${\bm{x}}_{k}={\bm{x}}_{k-1,T+1}$
  return  ${\bm{x}}_{K}$
3.1 Smooth Convex Setting
We first study the setting where the loss function of each task is convex and smooth, in contrast to prior works only focusing on nonsmooth settings (Bertsekas, 2011; Li et al., 2019, 2020). The main reasons we consider this case separately are that we can show faster convergence assuming smoothness and because it covers many regression tasks studied in prior CL works; see e.g., Evron et al. (2022, 2024). Note that with smoothness the proximal iteration is equivalent to the backward gradient step, i.e.,
${\bm{x}}_{k-1,t+1}={\bm{x}}_{k-1,t}-\eta_{k}\nabla f_{t}({\bm{x}}_{k-1,t+1}).$
Hence, much of the analysis from Section 2 can be adapted here, and the main difference lies in bounding the gap $f({\bm{x}}_{k})-f({\bm{z}}_{k-1})$ within each epoch as follows.
Lemma 3.1.
Under Assumptions 1 and 3, for any ${\bm{z}}\in\mathbb{R}^{d}$ that is fixed in the $k$-th cycle of Algorithm 2 and for $\alpha>0,\beta>0$ such that $\frac{1}{\alpha}+\frac{1}{\beta}\leq\frac{1}{2}$, if the step sizes satisfy $\eta_{k}\leq\frac{1}{\sqrt{\beta}TL}$, then we have for $k\in[K]$
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}\leq\;$ $\eta_{k}^{2}L\sum_{t=1}^{T-1}\Big{\|}\sum_{s=t+1}^{T}\nabla f_{s}({\bm{x}}_{*})\Big{\|}^{2}+\frac{\alpha}{\beta}T\big{(}f({\bm{z}})-f({\bm{x}}_{*})\big{)}$ (3.1)
$+\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}\|^{2}\big{)}.$
Proof.
Since each $f_{t}$ is convex and $L$-smooth, we have for $t\in[T]$
$f_{t}({\bm{x}}_{k})-f_{t}({\bm{x}}_{k-1,t+1})\leq\left\langle\nabla f_{t}({\bm{x}}_{k}),{\bm{x}}_{k}-{\bm{x}}_{k-1,t+1}\right\rangle-\frac{\|\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t+1})\|^{2}}{2L},$
$f_{t}({\bm{x}}_{k-1,t+1})-f_{t}({\bm{z}})\leq\left\langle\nabla f_{t}({\bm{x}}_{k-1,t+1}),{\bm{x}}_{k-1,t+1}-{\bm{z}}\right\rangle-\frac{\|\nabla f_{t}({\bm{x}}_{k-1,t+1})-\nabla f_{t}({\bm{z}})\|^{2}}{2L}.$
Following the proof of Lemma 2.1, we add and subtract $\frac{1}{2\eta_{k}}\|{\bm{x}}_{k-1,t}-{\bm{z}}\|^{2}$ on the right-hand side of the second inequality and combine the above two inequalities to obtain
$f_{t}({\bm{x}}_{k})-f_{t}({\bm{z}})\leq\;$ $\left\langle\nabla f_{t}({\bm{x}}_{k}),{\bm{x}}_{k}-{\bm{x}}_{k-1,t+1}\right\rangle-\frac{1}{2\eta_{k}}\|{\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1,t}\|^{2}$
$-\frac{1}{2L}\Big{(}\|\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t+1})\|^{2}+\|\nabla f_{t}({\bm{x}}_{k-1,t+1})-\nabla f_{t}({\bm{z}})\|^{2}\Big{)}$
$+\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1,t}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k-1,t+1}-{\bm{z}}\|^{2}\big{)}.$
Decomposing $\nabla f_{t}({\bm{x}}_{k})=\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t+1})+\nabla f_{t}({\bm{x}}_{k-1,t+1})$ and summing over $t\in[T]$, we note that ${\bm{x}}_{k-1}={\bm{x}}_{k-1,1}$ and ${\bm{x}}_{k}={\bm{x}}_{k-1,T+1}$ and obtain
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}$
[…]
$+\underbrace{\sum_{t=1}^{T}\left\langle\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t+1}),{\bm{x}}_{k}-{\bm{x}}_{k-1,t+1}\right\rangle}_{{\mathcal{T}}_{2}}+\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}\|^{2}\big{)}$
$-\frac{1}{2L}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t+1})\|^{2}-\frac{1}{2L}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{x}}_{k-1,t+1})-\nabla f_{t}({\bm{z}})\|^{2}.$
For the term ${\mathcal{T}}_{1}$, we follow the argument from the proof of Theorem 2.3 to obtain
${\mathcal{T}}_{1}=-\frac{1}{2\eta_{k}}\|{\bm{x}}_{k}-{\bm{x}}_{k-1}\|^{2}\leq 0.$
For the term ${\mathcal{T}}_{2}$, noticing that ${\bm{x}}_{k}-{\bm{x}}_{k-1,t+1}=-\eta_{k}\sum_{s=t+1}^{T}\nabla f_{s}({\bm{x}}_{k-1,s+1})$ for $11\leq t\leq T-$ 1 and decomposing $\nabla f_{s}({\bm{x}}_{k-1,s+1})=(\nabla f_{s}({\bm{x}}_{k-1,s+1})-\nabla f_{s}({\bm{z}}))+(\nabla f_{s}({\bm{z}})-\nabla f_{s}({\bm{x}}_{*}))+\nabla f_{s}({\bm{x}}_{*})$, we use Young’s inequality with parameters $\alpha>0$ and $\beta>0$ to obtain
${\mathcal{T}}_{2}=\;$ $\sum_{t=1}^{T-1}\Big{\langle}\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t+1}),-\eta_{k}\sum_{s=t+1}^{T}\nabla f_{s}({\bm{x}}_{k-1,s+1})\Big{\rangle}$
[…]
$+\frac{\alpha\eta_{k}^{2}L}{2}\sum_{t=1}^{T-1}\Big{\|}\sum_{s=t+1}^{T}\big{(}\nabla f_{s}({\bm{z}})-\nabla f_{s}({\bm{x}}_{*})\big{)}\Big{\|}^{2}+\eta_{k}^{2}L\sum_{t=1}^{T-1}\Big{\|}\sum_{s=t+1}^{T}\nabla f_{s}({\bm{x}}_{*})\Big{\|}^{2}$
$+\frac{\beta\eta_{k}^{2}L}{2}\sum_{t=1}^{T-1}\Big{\|}\sum_{s=t+1}^{T}\big{(}\nabla f_{s}({\bm{x}}_{k-1,s+1})-\nabla f_{s}({\bm{z}})\big{)}\Big{\|}^{2}.$
Further using the fact that $\|\sum_{i=1}^{n}{\bm{x}}_{i}\|^{2}\leq n\sum_{i=1}^{n}\|{\bm{x}}_{i}\|^{2}$ and combining the above bounds on ${\mathcal{T}}_{1}$ and ${\mathcal{T}}_{2}$, we obtain
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}$ $\leq\frac{1}{2L}\Big{(}\frac{1}{\alpha}+\frac{1}{\beta}-\frac{1}{2}\Big{)}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{x}}_{k})-\nabla f_{t}({\bm{x}}_{k-1,t})\|^{2}$
$\;+\Big{(}\frac{\beta\eta_{k}^{2}T^{2}L}{2}-\frac{1}{2L}\Big{)}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{x}}_{k-1,t+1})-\nabla f_{t}({\bm{z}})\|^{2}$
$\;+\frac{\alpha\eta_{k}^{2}T^{2}L}{2}\sum_{t=1}^{T}\|\nabla f_{t}({\bm{z}})-\nabla f_{t}({\bm{x}}_{*})\|^{2}+\eta_{k}^{2}L\sum_{t=1}^{T-1}\Big{\|}\sum_{s=t+1}^{T}\nabla f_{s}({\bm{x}}_{*})\Big{\|}^{2}$
$\;+\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}\|^{2}\big{)}.$
The rest of the proof is the same as the proof of Lemma 2.1 and is thus omitted. ∎
Choosing ${\bm{z}}={\bm{z}}_{k-1}$ defined by (2.5) in Lemma 3.1 and handling the error terms w.r.t. $\sigma_{*}^{2}$ and $f({\bm{z}}_{k-1})-f({\bm{x}}_{*})$ as in Theorem 2.3, we can obtain the last iterate guarantee for incremental proximal methods with the same rate, as stated in the following theorem.
Theorem 3.2 (IP).
Under Assumptions 13 and 4 and for parameters $\alpha,\beta$ such that $\frac{1}{\alpha}+\frac{1}{\beta}\leq\frac{1}{2}$, if the constant step size $\eta\leq\frac{1}{\sqrt{\beta}TL}$, the output ${\bm{x}}_{K}$ of Algorithm 2 satisfies
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\mathrm{e}\eta^{2}T^{2}\sigma_{*}^{2}L(1+\beta/\alpha)K^{\frac{\alpha/\beta}{1+\alpha/\beta}}+\frac{\mathrm{e}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{2T\eta K^{\frac{1}{1+\alpha/\beta}}}.$
With $\alpha=4,\beta=4\log K$, there exists a constant step size $\eta$ such that $f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\epsilon$ for $\epsilon>0$ after $\widetilde{\mathcal{O}}\big{(}\frac{TL\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon}+\frac{TL^{1/2}\sigma_{*}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{3/2}}\big{)}$ individual gradient evaluations.
Proof.
Follows from the proof of Theorem 2.3. ∎
A few remarks are in order here. First, the extensions to the increasing weighted averaging and RR/SO shuffling discussed in Section 2 also apply to this setting, which we omit for brevity. Second, in order to take advantage of smoothness to handle incremental updates, we require that the step size parameter $\eta$ is at most order-$1/(TL)$, which leads to less forgetting but larger regularization error. Similar step size restrictions also appear in prior work on incremental proximal method (Bertsekas, 2011; Li et al., 2019), requiring step sizes of the order $1/\sqrt{K}$ for convex Lipschitz components (same as our choice in the next subsection). On the other hand, weak regularization has only been studied asymptotically in Evron et al. (2023) for linear classification tasks with the regularization parameter tending to zero, in the need of relating to the sequential max-margin projections for their analysis. Evron et al. (2023) also left a concluding remark that another challenging but rewarding avenue is to extend our analysis to finite regularization strengths. Hence, at the current stage, it is unclear whether such a restriction on our small step size can be lifted. An interesting aspect of our result is that there is a critical value
$\eta_{*}=\min\Big{\{}\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2/3}}{2^{1/3}T\sigma_{*}^{2/3}L^{1/3}K^{1/3}(1+\beta/\alpha)^{1/3}},\,\frac{1}{\sqrt{\beta}TL}\Big{\}}$
such that if decrease $\eta$ beyond $\eta_{*}$, both the regularization error and the forgetting increase. It is thus an interesting open question to characterize how the forgetting is affected for $\eta<\eta_{*}$ (i.e., when we decrease the regularization error). Lastly, we note that for $\ell_{2}$-regularized CL, the role of the regularization parameter in the trade-off between the forgetting and the regularization error has only been theoretically calibrated for the setting of two linear regression tasks without cyclic replay (Heckel, 2022; Li et al., 2023), while the extensions to the cyclic replays or the general convex tasks remain unknown.
3.2 Convex Lipschitz Setting
We now further relax the smoothness assumption and consider the convex Lipschitz setting, with applications such as linear classification tasks considered in Evron et al. (2023).
To carry out the analysis, we leverage the standard fact that the proximal iteration is equivalent to the gradient step w.r.t. the Moreau envelope, i.e.,
${\bm{x}}_{k-1,t+1}={\bm{x}}_{k-1,t}-\eta_{k}\nabla M_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t}),$
while the gradient of the Moreau envelope $\nabla M_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})$ belongs to the subdifferential $\partial f_{t}({\bm{x}}_{k-1,t+1})$, thus is bounded by the Lipschitz constant of $f_{t}$. We use these observations to derive an upper bound on $f({\bm{x}}_{k})-f({\bm{z}}_{k-1})$ for each epoch, as summarized in the following lemma.
Lemma 3.3.
Under Assumptions 1 and 2, for any ${\bm{z}}\in\mathbb{R}^{d}$ that is fixed in the $k$-th cycle of Algorithm 2, we have for $k\in[K]$
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}\leq\;$ $\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}\|^{2}\big{)}+\frac{T(T-1)G^{2}\eta_{k}}{2}.$ (3.2)
Proof.
Since $f_{t}$ is convex and closed, we have
$\nabla M_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})=\frac{1}{\eta_{k}}({\bm{x}}_{k-1,t}-{\bm{x}}_{k-1,t+1})\in\partial f_{t}({\bm{x}}_{k-1,t+1}).$
By $G$-Lipschitzness of each component function, we have that for $t\in[T-1]$
$f_{t}({\bm{x}}_{k})-f_{t}({\bm{x}}_{k-1,t+1})$ (3.3)
[…]
On the other hand, using convexity of $f_{t}$, we have that for $t\in[T]$
$f_{t}({\bm{z}})\geq f_{t}({\bm{x}}_{k-1,t+1})+\left\langle\nabla M_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t}),{\bm{z}}-{\bm{x}}_{k-1,t+1}\right\rangle.$
Expanding the inner product in the above inequality leads to
$f_{t}({\bm{x}}_{k-1,t+1})-f_{t}({\bm{z}})$
[…]
=\;= $\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1,t}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k-1,t+1}-{\bm{z}}\|^{2}\big{)}-\frac{1}{2\eta_{k}}\|{\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1,t}\|^{2}.$ (3.4)
Combining (3.3) and (3.4) and noticing that ${\bm{x}}_{k-1,T+1}={\bm{x}}_{k}$ and ${\bm{x}}_{k-1,1}={\bm{x}}_{k-1}$, we sum the inequalities over $t\in[T]$ and obtain
$T(f({\bm{x}}_{k})-f({\bm{z}}))\leq\;$ $\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}\|^{2}\big{)}+\frac{T(T-1)G^{2}\eta_{k}}{2}$
$-\frac{1}{2\eta_{k}}\sum_{t=1}^{T}\|{\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1,t}\|^{2}$
[…]
We then use the sequence $\{{\bm{z}}_{k}\}$ defined in (2.5) to deduce the last iterate rate in the following theorem. In contrast to (2.1) and (3.1), we do not have the additional error term $f({\bm{z}})-f({\bm{x}}_{*})$ on the right-hand side of (3.2), so the analysis is much simplified.
Theorem 3.4 (IP).
Under Assumptions 1 and 2, the output ${\bm{x}}_{K}$ of Algorithm 2 satisfies
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\frac{1}{2T\sum_{k=1}^{K}\eta_{k}}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}+\frac{G^{2}T}{2}\sum_{k=1}^{K}\frac{\eta_{k}^{2}}{\sum_{j=k}^{K}\eta_{j}}.$
Moreover, given $\epsilon>0$, there exists a constant step size $\eta$ such that $f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\epsilon$ after $\widetilde{\mathcal{O}}\big{(}\frac{G^{2}T\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{2}}\big{)}$ individual gradient evaluations.
Proof.
Plugging ${\bm{z}}_{k-1}$ defined in (2.5) into (3.2) and multiplying $\eta_{k}w_{k-1}$ on both sides, we obtain
$T\eta_{k}w_{k-1}\big{(}f({\bm{x}}_{k})-f({\bm{z}}_{k-1})\big{)}$
[…]
Summing over $k\in[K]$ and using the second part of Lemma 2.2, we have
$\sum_{k=1}^{K}T\eta_{k}\Big{[}w_{k-1}(f({\bm{x}}_{k})-f({\bm{x}}_{*}))-\sum_{j=0}^{k-1}w_{j}(1-\lambda_{j})(f({\bm{x}}_{j})-f({\bm{x}}_{*}))\Big{]}$
[…]
where we also recall that ${\bm{z}}_{-1}={\bm{x}}_{*}$. Unrolling the terms on the left-hand side as (2.9) and choosing $\lambda_{0}=1$, we obtain
$T\eta_{K}w_{K-1}\big{(}f({\bm{x}}_{K})-f({\bm{x}}_{*})\big{)}$ (3.5)
$+T\sum_{k=1}^{K-1}\Big{[}\eta_{k}w_{k-1}-w_{k}(1-\lambda_{k})\sum_{j=k+1}^{K}\eta_{j}\Big{]}\big{(}f({\bm{x}}_{k})-f({\bm{x}}_{*})\big{)}$
[…]
To obtain the last iterate guarantee, we choose $\lambda_{k}$ and $w_{k}$ such that
$\lambda_{k}w_{k}\leq\;$ $w_{k-1},\quad 0\leq k\leq K-1,$
$\eta_{k}w_{k-1}-w_{k}(1-\lambda_{k})\sum_{j=k+1}^{K}\eta_{j}\geq\;$ $1\displaystyle 0,\quad 1\leq k\leq K-$ 1 .
For simplicity and without loss of generality, we make both inequalities tight and choose $w_{k}=\frac{\sum_{j=k}^{K}\eta_{j}}{\sum_{j=k+1}^{K}\eta_{j}}w_{k-1}$. In particular, we choose $w_{k}=\frac{\eta_{K}}{\sum_{j=k+1}^{K}\eta_{j}}$ for $10\leq k\leq K-$ 1 such that $w_{K-1}=1$, then we divide $T\eta_{K}$ on both sides of (3.5) and obtain
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\frac{w_{-1}}{2T\eta_{K}}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}+\frac{G^{2}T}{2\eta_{K}}\sum_{k=1}^{K}\eta_{k}^{2}w_{k-1}$
=\;= $\frac{1}{2T\sum_{k=1}^{K}\eta_{k}}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}+\frac{G^{2}T}{2}\sum_{k=1}^{K}\frac{\eta_{k}^{2}}{\sum_{j=k}^{K}\eta_{j}}.$
Finally, choosing $\eta_{k}\equiv\eta=\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{GT\sqrt{K}}$, we get
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\frac{G\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{2\sqrt{K}}\Big{(}1+\sum_{k=1}^{K}\frac{1}{K-k+1}\Big{)}\leq\frac{G\|{\bm{x}}_{0}-{\bm{x}}_{*}\|(1+\log K/2)}{\sqrt{K}}.$
Hence, given $\epsilon>0$, to guarantee $f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\epsilon$, the total number of individual gradient evaluations will be
$TK=\widetilde{\mathcal{O}}\Big{(}\frac{G^{2}T\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{2}}\Big{)},$
completing the proof. ∎
We note that the last iterate rate we obtained in Theorem 3.4 matches the prior best known results for the average iterate guarantees for incremental proximal methods (Bertsekas, 2011; Li et al., 2019) in nonsmooth settings, up to a logarithmic factor. Further, we take the $\Theta(\frac{1}{\sqrt{K}})$ step size in the proof only for analytical simplicity, while the diminishing step sizes $\eta_{k}=\Theta(\frac{1}{\sqrt{k}})$ will yield the same rate via a similar analysis, which we omit for brevity.
3.3 Inexact Proximal Point Evaluations
In the last two subsections, we derived our results assuming that the proximal point operator can be evaluated exactly. However, computing the proximal point corresponds to solving a strongly convex problem, which is generally possible to do only up to finite accuracy. When the component functions are smooth, the subproblems resulting from the proximal point evaluations are both smooth and strongly convex and as such can be solved efficiently to high accuracy (in ${\mathcal{O}}(\log(1/\epsilon))$ iterations to reach an error $\epsilon$). For brevity, we thus focus on the nonsmooth convex Lipschitz setting and discuss how to extend our results to admit certain inexactness arisen from approximating the proximal point. In particular, we use the inexactness assumption from Lemaire (1992); Cominetti (1997); Alber et al. (1997); Salzo and Villa (2012); Villa et al. (2013), i.e., where ${\bm{x}}_{k-1,t+1}$ is an approximation of $\mathrm{prox}_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})$ with error $\varepsilon_{k-1,t}>0$ if
$\frac{1}{\eta_{k}}({\bm{x}}_{k-1,t}-{\bm{x}}_{k-1,t+1})\in\partial_{\frac{\varepsilon_{k-1,t}^{2}}{2\eta_{k}}}f_{t}({\bm{x}}_{k-1,t+1}).$ (3.6)
By the results in Salzo and Villa (2012, Proposition 1), the inexact proximal point ${\bm{x}}_{k-1,t+1}$ also satisfies $f_{t}({\bm{x}}_{k-1,t+1})+\frac{1}{2\eta_{k}}\|{\bm{x}}_{k-1,t+1}-{\bm{x}}_{k-1,t}\|^{2}\leq\min_{{\bm{x}}\in\mathbb{R}^{d}}\Big{\{}f_{t}({\bm{x}})+\frac{1}{2\eta_{k}}\|{\bm{x}}-{\bm{x}}_{k-1,t}\|^{2}\Big{\}}+\frac{\varepsilon_{k-1,t}^{2}}{2\eta_{k}}$. Further using the strong convexity of $f_{t}({\bm{x}})+\frac{1}{2\eta_{k}}\|{\bm{x}}-{\bm{x}}_{k-1,t}\|^{2}$, we have
$\|{\bm{x}}_{k-1,t+1}-\mathrm{prox}_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})\|\leq\varepsilon_{k-1,t}.$
Denoting ${\bm{g}}_{k-1,t}:=\frac{1}{\eta_{k}}({\bm{x}}_{k-1,t}-{\bm{x}}_{k-1,t+1})$ and noticing that $\nabla M_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})=\frac{1}{\eta_{k}}({\bm{x}}_{k-1,t}-\mathrm{prox}_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t}))$, we have
$\|{\bm{g}}_{k-1,t}-\nabla M_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})\|\leq\frac{\varepsilon_{k-1,t}}{\eta_{k}}.$ (3.7)
With these insights, we are ready to bound $f({\bm{x}}_{k})-f({\bm{z}}_{k-1})$ as follows.
Lemma 3.5.
Under Assumptions 1 and 2, for any ${\bm{z}}\in\mathbb{R}^{d}$ that is fixed in the $k$-th cycle of Algorithm 2, we have for $k\in[K]$
$T\big{(}f({\bm{x}}_{k})-f({\bm{z}})\big{)}\leq\;$ $\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}\|^{2}\big{)}+\frac{T(T-1)G^{2}\eta_{k}}{2}$ (3.8)
$+\frac{1}{2\eta_{k}}\sum_{t=1}^{T}\varepsilon_{k-1,t}^{2}+GT\sum_{t=1}^{T}\varepsilon_{k-1,t}.$
Proof.
Since $f_{t}$ is convex and closed, we have
$\nabla M_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})=\frac{1}{\eta_{k}}({\bm{x}}_{k-1,t}-\mathrm{prox}_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t}))\in\partial f_{t}(\mathrm{prox}_{\eta_{k}f_{t}}({\bm{x}}_{k-1,t})).$
By Lipschitzness of each component function, we have for $t\in[T-1]$
$f_{t}({\bm{x}}_{k})-f_{t}({\bm{x}}_{k-1,t+1})\leq\;$ $G\|{\bm{x}}_{k}-{\bm{x}}_{k-1,t+1}\|=G\eta_{k}\Big{\|}\sum_{s=t+1}^{T}{\bm{g}}_{k-1,s}\Big{\|}.$
Decomposing ${\bm{g}}_{k-1,s}={\bm{g}}_{k-1,s}-\nabla M_{\eta_{k}f_{s}}({\bm{x}}_{k-1,s})+\nabla M_{\eta_{k}f_{s}}({\bm{x}}_{k-1,s})$ and using triangle inequality, we have
$f_{t}({\bm{x}}_{k})-f_{t}({\bm{x}}_{k-1,t+1})$
[…]
$\overset{(i)}{\leq}\;$ $\eta_{k}G\sum_{s=t+1}^{T}\Big{(}\frac{\varepsilon_{k-1,s}}{\eta_{k}}+G\Big{)}\leq(T-t)G^{2}\eta_{k}+G\sum_{s=t+1}^{T}\varepsilon_{k-1,s},$ (3.9)
where we use (3.7) and 2 for $(i)$. On the other hand, by the definition of $\frac{\epsilon_{k-1,t}^{2}}{2\eta_{k}}$-subdifferential of $f_{t}$ and using (3.6), we obtain
$f_{t}({\bm{z}})\geq f_{t}({\bm{x}}_{k-1,t+1})+\left\langle{\bm{g}}_{k-1,t},{\bm{z}}-{\bm{x}}_{k-1,t+1}\right\rangle-\frac{\varepsilon_{k-1,t}^{2}}{2\eta_{k}}.$
Expanding the inner product in the above quantity leads to
$f_{t}({\bm{x}}_{k-1,t+1})-f_{t}({\bm{z}})$
[…]
Combining (3.9) and (3.10) and noticing that ${\bm{x}}_{k-1,T+1}={\bm{x}}_{k}$ and ${\bm{x}}_{k-1,1}={\bm{x}}_{k-1}$, we sum the inequalities over $t\in[T]$ and obtain
$T(f({\bm{x}}_{k})-f({\bm{z}}))\leq\;$ $\frac{1}{2\eta_{k}}\big{(}\|{\bm{x}}_{k-1}-{\bm{z}}\|^{2}-\|{\bm{x}}_{k}-{\bm{z}}\|^{2}\big{)}+\frac{T(T-1)G^{2}\eta_{k}}{2}$
$+\frac{1}{2\eta_{k}}\sum_{t=1}^{T}\varepsilon_{k-1,t}^{2}+GT\sum_{t=1}^{T}\varepsilon_{k-1,t}.$
Compared to (3.2) in Lemma 3.3, we now have additional error terms in (3.8) from the inexact evaluations. However, as long as the approximation error at each iteration is small, i.e., solving the subproblem of proximal point evaluation to a certain accuracy, we can still maintain the convergence rate of incremental proximal methods with exact proximal point evaluations. We formalize this in the following theorem.
Theorem 3.6 (Inexact Evaluation).
Under Assumptions 1 and 2, the output ${\bm{x}}_{K}$ of Algorithm 2 satisfies
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\frac{1}{2T\sum_{k=1}^{K}\eta_{k}}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}+\frac{G^{2}T}{2}\sum_{k=1}^{K}\frac{\eta_{k}^{2}}{\sum_{j=k}^{K}\eta_{j}}$
$+\frac{1}{2T}\sum_{k=1}^{K}\sum_{t=1}^{T}\frac{\varepsilon_{k-1,t}^{2}}{\sum_{j=k}^{K}\eta_{j}}+G\sum_{k=1}^{K}\sum_{t=1}^{T}\frac{\varepsilon_{k-1,t}\eta_{k}}{\sum_{j=k}^{K}\eta_{j}}.$
Further, given $\epsilon>0$, if $\sum_{t=1}^{T}\varepsilon_{k-1,t}=\tilde{\mathcal{O}}\big{(}\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{\sqrt{K}}\big{)}$, then there exists a constant step size $\eta$ such that $f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\epsilon$ after $\widetilde{\mathcal{O}}\big{(}\frac{G^{2}T\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{2}}\big{)}$ individual gradient evaluations.
Proof.
Using Lemma 3.5 and choosing ${\bm{z}}={\bm{z}}_{k-1}$ defined by (2.5) and following the proof of Theorem 3.4 with $w_{k}=\frac{\eta_{K}}{\sum_{j=k+1}^{K}\eta_{j}}$, we arrive at
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\frac{w_{-1}}{2T\eta_{K}}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}+\frac{G^{2}T}{2\eta_{K}}\sum_{k=1}^{K}\eta_{k}^{2}w_{k-1}$
$+\frac{1}{2T\eta_{K}}\sum_{k=1}^{K}\sum_{t=1}^{T}\varepsilon_{k-1,t}^{2}w_{k-1}+\frac{G}{\eta_{K}}\sum_{k=1}^{K}\sum_{t=1}^{T}\varepsilon_{k-1,t}\eta_{k}w_{k-1}$
=\;= $\frac{1}{2T\sum_{k=1}^{K}\eta_{k}}\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}+\frac{G^{2}T}{2}\sum_{k=1}^{K}\frac{\eta_{k}^{2}}{\sum_{j=k}^{K}\eta_{j}}$
$+\frac{1}{2T}\sum_{k=1}^{K}\sum_{t=1}^{T}\frac{\varepsilon_{k-1,t}^{2}}{\sum_{j=k}^{K}\eta_{j}}+G\sum_{k=1}^{K}\sum_{t=1}^{T}\frac{\varepsilon_{k-1,t}\eta_{k}}{\sum_{j=k}^{K}\eta_{j}}.$
To analyze the oracle complexity, we choose the constant step size $\eta_{k}\equiv\eta=\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{GT\sqrt{K}}$ for simplicity and without loss generality, which leads to
$f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\;$ $\frac{G\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{2\sqrt{K}}\Big{(}1+\sum_{k=1}^{K}\frac{1}{K-k+1}\Big{)}$
$+\frac{G\sqrt{K}}{2\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}\sum_{k=1}^{K}\frac{\sum_{t=1}^{T}\varepsilon_{k-1,t}^{2}}{K-k+1}+G\sum_{k=1}^{K}\frac{\sum_{t=1}^{T}\varepsilon_{k-1,t}}{K-k+1}.$
If $\sum_{t=1}^{T}\varepsilon_{k-1,t}=\tilde{\mathcal{O}}\big{(}\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{\sqrt{K}}\big{)}$ and thus $\sum_{t=1}^{T}\varepsilon_{k-1,t}^{2}=\tilde{\mathcal{O}}\big{(}\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{K}\big{)}$ (for example letting $\varepsilon_{k-1,t}=\frac{\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{T\sqrt{K}}$), we can then guarantee $f({\bm{x}}_{K})-f({\bm{x}}_{*})=\tilde{\mathcal{O}}\Big{(}\frac{G\|{\bm{x}}_{0}-{\bm{x}}_{*}\|}{\sqrt{K}}\Big{)}$. Hence, given $\epsilon>0$, to guarantee $f({\bm{x}}_{K})-f({\bm{x}}_{*})\leq\epsilon$, the total number of individual gradient evaluations will be
$KT=\tilde{\mathcal{O}}\Big{(}\frac{G^{2}T\|{\bm{x}}_{0}-{\bm{x}}_{*}\|^{2}}{\epsilon^{2}}\Big{)}.$
4 Conclusion
This work provides the first oracle complexity guarantees for the last iterate of standard incremental gradient methods, motivated by applications in continual learning. The obtained complexity bounds nearly match the best known oracle complexity bounds that in the same settings were previously known only for the (uniformly) average iterate. We further provided generalizations of these results to weighted average iterate with increasing weights and variants of methods with permuted ordering of updates. Some questions that merit further investigation remain. In particular, our results for the proximal incremental method – which are most closely related to continual learning settings – require a large value of the regularization parameter to bound the forgetting. Further increasing regularization deteriorates both the bounds on the forgetting and the loss on the current task, so this value in our results can be seen as a “critical value.” However, how the forgetting is affected if the regularization is reduced (which improves the loss on the current task) is not known. This question is also closely related to whether larger theoretically-justified step sizes are possible for incremental gradient/shuffled SGD methods, which has remained open despite renewed interest and advances in theory of shuffled SGD convergence.
Acknowledgments
This research was supported in part by the U.S. Office of Naval Research under contract number N00014-22-1-2348.
Supplementary Material
Appendix A Omitted Proofs
See 2.6
Proof.
We first consider the case of random reshuffling strategy. Conditional on all the randomness up to but not including $k$-th epoch, the only randomness of $\mathbb{E}_{k}[\sum_{t=1}^{T}\|\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{*})\|^{2}]$ comes from the permutation $\pi^{(k)}$ at $k$-th epoch. Further noticing that each partial sum $\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{*})$ can be seen as a batch sampled without replacement from $\{\nabla f_{t}({\bm{x}}_{*})\}_{t\in[T]}$, we have
$\mathbb{E}_{k}\Big{[}\sum_{t=1}^{T}\Big{\|}\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{*})\Big{\|}^{2}\Big{]}=\;$ $\sum_{t=1}^{T}\mathbb{E}_{\pi^{(k)}}\Big{[}\Big{\|}\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{*})\Big{\|}^{2}\Big{]}$
=\;= $\sum_{t=1}^{T}(T-t+1)^{2}\mathbb{E}_{\pi^{(k)}}\Big{[}\Big{\|}\frac{1}{T-t+1}\sum_{s=t}^{T}\nabla f_{s}({\bm{x}}_{*})\Big{\|}^{2}\Big{]}$
$\overset{(i)}{\leq}\;$ $\sum_{t=1}^{T}(T-t+1)^{2}\frac{t-1}{(T-t+1)(T-1)}\sigma_{*}^{2}$
=\;= $\frac{T(T+1)}{6}\sigma_{*}^{2},$
where $(i)$ is due to $\sum_{t=1}^{T}\nabla f_{t}({\bm{x}}_{*})=\mathbf{0}$ and sampling without replacement (see e.g., (Lohr, 2021, Section 2.7)). It remains to take expectation w.r.t. all randomness on both sides and use the law of total expectation. For the case of shuffle-once variant, we can directly take expectation since the randomness only comes from the initial random permutation, and the above argument still applies. ∎